# PlaceFormer: Transformer-based Visual Place Recognition using Multi-Scale Patch Selection and Fusion

### v1

@@ I Introduction
Visual Place Recognition (VPR) is a critical task for localizing autonomous vehicles and robots navigating through dynamic environments, relying on visual input such as images or videos. Visual place recognition is defined as an image retrieval problem [1], where a query image from an unknown location is compared with a database of reference images from known locations in order to localize the query image. The location of the query image is estimated by identifying the closest matching image in the reference image database. This task is challenging due to variations in seasons, illumination, viewpoint, and occlusions. Typically, two types of image representations are used in VPR tasks: global and patch-level descriptors. Global descriptors [2, 3, 4] provide a succinct image representation in a single vector, facilitating efficient large-scale searches. Patch-level or local descriptors [5, 6, 7] encode details about specific regions or key points of the image and are used for performing geometric verification between image pairs.
@@
In this paper, we propose PlaceFormer, a novel approach that harnesses the potential of vision transformers to extract robust image representations specifically designed for visual place recognition. The global retrieval process involves aggregating the patch tokens of the transformer and utilizing the same to perform global retrieval. Then, we compute patches of multiple fixed scales on the images by fusing the patch tokens and then, identify key patches amongst them, leveraging the attention map of the vision transformer. These key patches pinpoint areas of the image ideal for accurate long-term VPR. Comparing key patches in both query and reference images across different scales, we compute similarity scores using geometric verification. Subsequently, a re-ranking process is carried out based on these similarity scores. In Fig. 1, we illustrate a visual example that highlights key patches of various scales selected based on attention scores (marked with boxes of distinct colors). The green lines represent the correspondences estimated between patches of different scales in both the query and reference images, which are utilized for the re-ranking process.
@@ II Related Works
Patch-Level Descriptors. Earlier approaches for extracting patch-level descriptors relied on handcrafted features such as SIFT [15], SURF [16], and BRIEF [17] at key points. However, these features struggled to adapt to the substantial long-term changes typical in place recognition tasks. CNNs have also been employed for patch-level descriptor extraction [18, 19, 13, 7, 20], capturing features from diverse image regions. Patch-NetVLAD [6] adapted the NetVLAD global descriptor framework to create descriptors for multiple fixed-size patches in an image. Expanding on this idea, Hot-NetVLAD [5] proposes techniques to identify image patches crucial for place recognition. Focusing on these specific areas optimizes Patch-NetVLAD’s application, reducing memory usage and computational demands compared to methods processing all available patches.
TransVPR [21] combines CNN and vision transformer for place recognition, extracting global and local features by integrating a CNN backbone with transformer layers. Particularly noteworthy is its selection of vital local features for place recognition through the merging of attention maps from diverse transformer layers. $R^{2}$ormer [22], another transformer-based method, extends beyond mere feature extraction, employing transformers for both the feature extraction and re-ranking process. These transformer-based methodologies rely on the use of patch tokens from transformers as such to extract patch descriptors. However, the direct utilization of patch tokens may not always yield comprehensive correspondences when matching two images, especially when there are variations between the images. To overcome this, we propose a local fusion of patch tokens to extract patches of various scales and cross-matching patches of different scales, hence estimating extensive correspondences between the images for the matching process.
@@ III Methodology
Figure 2: The proposed framework, PlaceFormer, encompasses a two-phase approach for visual place recognition. In the global retrieval phase, patch tokens extracted from the vision transformer undergo pooling and are subsequently processed through a linear layer, resulting in a feature vector utilized for efficient global retrieval. In the re-ranking phase, the patch tokens and attention map from the transformer’s last layer are fused to generate patches at multiple scales. Leveraging attention scores, key patches are selectively identified, and correspondences between patches of different scales are computed. In the figure, for brevity, only a few inliers between the patches have been visualized. These inliers contribute to computing a spatial matching score, crucial for the re-ranking process.
@@ III-C Attention-based Key Patch Selection
To optimize the re-ranking of candidate images and minimize computational overhead, we concentrate on key patch tokens identified through the attention map of the transformer. This strategy emphasizes the most pertinent patches, improving efficiency without compromising accuracy. Utilizing the attention maps $A_{L}$, $A_{2L}$, and $A_{3L}$, we choose the top $400$, $200$, and $50$ patches, respectively, from $P_{L}$, $P_{L2}$, and $P_{L3}$ based on their attention scores. Subsequently, we refine the selection by filtering the chosen patch tokens, retaining only those with an attention score surpassing a predefined threshold $\tau$, consequently identifying the key patches across the three scales $P_{K}$, $P_{K2}$, and $P_{K3}$, where $P_{K}\subseteq P_{L}$, $P_{K2}\subseteq P_{L2}$ and $P_{K3}\subseteq P_{L3}$.
@@ III-E Multi-Scale Patch Matching
The final spatial score $S_{spatial}$ is computed by the summation of the spatial matching scores computed across patches of different scales,
@@ IV-A Implementation and Training Details
PlaceFormer is developed using the PyTorch framework [24]. The training and testing of the models are performed on an NVIDIA RTX 3090 graphics card. For the base encoder, we utilize the Vision Transformer Small (ViT-S) model [8]. This model is characterized by its $12$ layers of transformers, each containing $12$ heads. The transformer within this model is designed to extract patch tokens, each with a dimension ($d$) of $384$. We chose patch size, $p$ of $16\times 16$ aligning with the model’s architecture. To maintain consistency with previous research and ensure compatibility, all images used in both training and testing phases are resized to a resolution of $640\times 480$. A key patch filtering threshold, $\tau$ of $0.01$ is used in the implementation. The re-ranking is performed on top-100 ($k=100$) candidates retrieved through global retrieval. A margin, $m$ of $0.01$ is used in the triplet loss.
Training. The transformer is initialized with pre-trained weights on ImageNet-1K for the training process. The model is trained using the MSLS train dataset [25]. MSLS is chosen as the train set due to its diversity in scene types and the presence of various environmental variations, providing a comprehensive training environment. The model is further fine-tuned using Pittsburgh 30K (Pitts30K) training set [26].
Both the positive and negative samples are pre-computed before training to optimize the training duration. The training utilized the Adam optimizer, in conjunction with a cosine learning rate scheduler. The initial learning rate is set at $0.0005$. The training is continued until there is no further improvement in accuracy on the validation set, ensuring that the model is fully optimized and capable of accurately recognizing and differentiating between a wide range of scenes and environmental conditions.
@@
Method MSLS Validation [25] MSLS Challenge [25] Nordland [27] Pitts30K [26] Tokyo 24/7 [28]
R@1
R@5
R@10 R@1
R@5
R@10 R@1
R@5
R@10 R@1
R@5
R@10 R@1
R@5
R@10
NetVLAD [2] 60.8
74.3
79.5 35.1
47.4
51.7 13.6
21.4
25.2 81.9
91.2
93.7 64.8
78.4
81.6
SFRS [29] 69.2
80.3
83.1 41.5
52.0
56.3 16.4
26.3
29.7 89.4
94.7
95.9 85.4
91.1
93.3
PlaceFormer (w/o re-ranking) 80.0
90.0
93.0 58.0
76.7
81.6 26.1
40.3
47.6 82.7
93.0
95.1 57.5
74.6
80.3
SP-SuperGlue [30] 78.1
81.9
84.3 50.6
56.9
58.3 29.1
33.5
34.3 87.2
94.8
96.4 88.2
90.2
90.2
DELG [13] 83.2
89.3
91.1 52.2
61.9
65.4 51.0
63.9
66.7 89.8
95.3
96.7 86.4 92.4 93.0
Patch-NetVLAD [6] 79.5
86.2
87.7 48.1
57.6
60.5 46.4
58.0
60.4 88.7
94.5
95.9 86.0
88.6
90.5
TransVPR [21] 86.8
91.2
92.4 63.9
74.0
77.5 58.8 75.0 78.7 89.0
94.9
96.2 79.0
82.2
85.1
$R^{2}$ormer [22] 89.7 95.0 96.2 73.0 85.9 88.8 60.6
66.8
68.7 91.1
95.2
96.3 88.6
91.4
91.7
PlaceFormer (Ours) 89.9
94.3
95.4 71.9
85.4
88.7 65.3
70.5
72.4 92.4 96.5 97.4 87.6
89.2
91.5
@@ IV-B Datasets
We evaluate PlaceFormer on multiple public benchmark datasets. These included the Nordland [27], Pittsburgh 30K (Pitts30K) [26], Tokyo24/7 [28] and Mapillary Street-Level Sequences (MSLS) [25]. Each of these datasets offers a unique set of challenges, encompassing a wide array of environments and conditions that are crucial for thorough performance assessment. All images used in the evaluation are resized to a uniform resolution of $640\times 480$ to be consistent with other methods. The model trained on MSLS is used to evaluate MSLS and Nordland datasets. The model fine-tuned on Pitts30K is used to evaluate Pitts30K and Tokyo 24/7 (urban scenarios).
@@ IV-C Metrics
Recall@K is used as the primary metric for evaluation. This metric quantifies the percentage of query images that are correctly localized within a dataset. It does so by determining whether at least one of the top K-ranked reference images falls within a specified threshold distance from the query image. For our evaluation, we followed the precedent set in prior works [23, 21, 6], using a threshold distance of $25$ meters. This distance is chosen as it represents a reasonable and commonly accepted proximity for accurate localization in urban environments. The Recall@K is measured for K values of 1, 5, and 10, providing a comprehensive understanding of the model’s performance across different levels of retrieval granularity.
@@ IV-D Comparison with State-of-the-arts
In our comparative analysis, PlaceFormer is benchmarked against a range of state-of-the-art methods to demonstrate its efficacy in visual place recognition. We evaluated it alongside methods such as NetVLAD [2] and SFRS [29], which primarily utilize global image representations. Alongside these methods, we also presented the results of our global retrieval to offer a comprehensive comparison. Furthermore, we compared PlaceFormer with techniques that employ both global and local features for retrieval and ranking. This included comparisons with Patch-NetVLAD [6] and DELG [13].
@@ V-A Quantitative Results
The quantitative performance of PlaceFormer, in comparison to other approaches, is detailed in Table I. PlaceFormer without local re-ranking outperforms other global retrieval methods in MSLS validation, MSLS challenge, and Nordland datasets, and yields comparable performance in Pitts30K and Tokyo 24/7 datasets. Furthermore, PlaceFormer without re-ranking outperforms multiple re-ranking methods as well, making it suitable to be used even without re-ranking based on specific requirements.
PlaceFormer with re-ranking achieves competitive results on all datasets. It outperforms all comparable methods on MSLS validation, Nordland, and Pitts30K in Recall@1 with absolute differences of $0.2\%$, $5.3\%$, and $1.3\%$ compared to the best-performing baseline of $R^{2}$ormer. When computing the average performance across all datasets, PlaceFormer demonstrates a substantial superiority over competing methods, surpassing NetVLAD, SFRS, SP-SuperGlue, DELG, PatchNetVLAD, TransVPR, and $R^{2}$ormer by margins of $30.8\%$, $21.04\%$, $14.78\%$, $8.9\%$, $11.68\%$, $5.92\%$, and $0.82\%$ in Recall@1 scores, respectively. While PlaceFormer generally surpasses other state-of-the-art methods, it falls short of outperforming $R^{2}$ormer on certain datasets such as MSLS Challenge and Tokyo 24/7. It is important to highlight that our method is specifically trained for global retrieval only, whereas $R^{2}$ormer undergoes training for re-ranking, contributing to its enhanced performance.

### v2

@@ I Introduction
Visual Place Recognition (VPR) is a critical task for localizing autonomous vehicles and robots navigating through dynamic environments, relying on visual input such as images or videos. Visual place recognition is defined as an image retrieval problem [1], where a query image from an unknown location is compared with a database of reference images from known locations in order to localize the query image. The location of the query image is estimated by identifying the closest matching image in the reference image database. This task is challenging due to variations in seasons, illumination, viewpoint, and occlusions. Typically, two types of image representations are used in VPR tasks: global and patch-level descriptors. Global descriptors [2, 3] provide a succinct image representation in a single vector, facilitating efficient large-scale searches. Patch-level or local descriptors [4, 5, 6] encode details about specific regions or key points of the image and are used for performing geometric verification between image pairs.
@@
In this paper, we propose PlaceFormer, a novel approach that harnesses the potential of vision transformers to extract robust image representations specifically designed for visual place recognition. The global retrieval process involves aggregating the patch tokens of the transformer and utilizing the same to perform global retrieval. Then, we compute patches of multiple fixed scales on the images by fusing the patch tokens and then, identify key patches amongst them, leveraging the attention map of the vision transformer. Patches of multiple scales are employed to enhance correspondence matching between images despite variations in the scale and viewpoints.. These key patches pinpoint areas of the image ideal for accurate long-term VPR. Comparing key patches in both query and reference images across different scales, we compute similarity scores using geometric verification. Subsequently, a re-ranking process is carried out based on these similarity scores. In Fig. 1, we illustrate a visual example that highlights key patches of various scales selected based on attention scores (marked with boxes of distinct colors). The green lines represent the correspondences estimated between patches of different scales in both the query and reference images, which are utilized for the re-ranking process.
@@ II Related Works
Patch-Level Descriptors. Earlier approaches for extracting patch-level descriptors relied on handcrafted features such as SIFT [14], SURF [15], and BRIEF [16] at key points. However, these features struggled to adapt to the substantial long-term changes typical in place recognition tasks. CNNs have also been employed for patch-level descriptor extraction [17, 18, 12, 6], capturing features from diverse image regions. Patch-NetVLAD [5] adapted the NetVLAD global descriptor framework to create descriptors for multiple fixed-size patches in an image. Expanding on this idea, Hot-NetVLAD [4] proposes techniques to identify image patches crucial for place recognition. Focusing on these specific areas optimizes Patch-NetVLAD’s application, reducing memory usage and computational demands compared to methods processing all available patches.
TransVPR [19] combines CNN and vision transformer for place recognition, extracting global and local features by integrating a CNN backbone with transformer layers. Particularly noteworthy is its selection of vital local features for place recognition through the merging of attention maps from diverse transformer layers. $R^{2}$ormer [20], another transformer-based method, extends beyond mere feature extraction, employing transformers for both the feature extraction and re-ranking process. These transformer-based methodologies rely on the use of patch tokens from transformers as such to extract patch descriptors. These techniques leverage patch tokens of the same size from vision transformers to estimate correspondences when matching two images. However, variations in viewpoint and scale across images may cause some correspondences to be overlooked, consequently impacting performance. To overcome this, we propose a local fusion of patch tokens to extract patches of various scales and cross-matching patches of different scales, hence estimating extensive correspondences between the images for the matching process.
@@ III Methodology
Figure 2: The proposed framework, PlaceFormer, encompasses a two-phase approach for visual place recognition. In the global retrieval phase, patch tokens extracted from the vision transformer undergo pooling and are subsequently processed through a linear layer, resulting in a feature vector utilized for efficient global retrieval. In the re-ranking phase, the patch tokens and attention map from the transformer’s last layer are fused to generate patches at multiple scales. Leveraging attention scores, key patches are selectively identified, and correspondences between patches of different scales are computed. In the figure, for brevity, only a few inliers between the patches have been visualized. These inliers contribute to computing a spatial matching score, which is crucial for the re-ranking process.
@@ III-C Attention-based Key Patch Selection
To optimize the re-ranking of candidate images and minimize computational overhead, we concentrate on key patch tokens identified through the attention map of the transformer. This strategy emphasizes the most pertinent patches, improving efficiency without compromising accuracy. Utilizing the attention maps $A_{L}$, $A_{2L}$, and $A_{3L}$, we choose the top $400$, $200$, and $50$ patches [^1] The limit on the number of patches was determined through experimentation in the development phase. Based on the attention score threshold, $\tau$, the chosen number of patches consistently ranged around 400, 200, and 50 for each respective patch size. These findings led to capping the maximum number of patches to establish an upper limit on memory requirements. It was also found that increasing these limits did not lead to any significant increase in performance., respectively, from $P_{L}$, $P_{L2}$, and $P_{L3}$ based on their attention scores. Subsequently, we refine the selection by filtering the chosen patch tokens, retaining only those with an attention score surpassing a predefined threshold $\tau$, consequently identifying the key patches across the three scales $P_{K}$, $P_{K2}$, and $P_{K3}$, where $P_{K}\subseteq P_{L}$, $P_{K2}\subseteq P_{L2}$ and $P_{K3}\subseteq P_{L3}$.
@@ III-E Multi-Scale Patch Matching
The final spatial score $S_{spatial}$ [^2] Weights were introduced during development to scale individual spatial matching scores, yet no significant changes in the results were observed. Consequently, these weights were omitted from the final method. is computed by the summation of the spatial matching scores computed across patches of different scales,
@@ IV-A Implementation and Training Details
PlaceFormer is developed using the PyTorch framework. The training and testing of the models are performed on an NVIDIA RTX 3090 graphics card. For the base encoder, we utilize the Vision Transformer Small (ViT-S) model [7]. This model is characterized by its $12$ layers of transformers, each containing $12$ heads. The transformer within this model is designed to extract patch tokens, each with a dimension ($d$) of $384$. We chose patch size, $p$ of $16\times 16$ aligning with the model’s architecture. To maintain consistency with previous research and ensure compatibility, all images used in both training and testing phases are resized to a resolution of $640\times 480$. A key patch filtering threshold, $\tau$ of $0.01$ is used in the implementation. The re-ranking is performed on top-100 ($k=100$) candidates retrieved through global retrieval. A margin, $m$ of $0.01$ is used in the triplet loss.
Training. The transformer is initialized with pre-trained weights on ImageNet-1K for the training process. The model is trained using the MSLS train dataset [22]. MSLS is chosen as the train set due to its diversity in scene types and the presence of various environmental variations, providing a comprehensive training environment. The model is further fine-tuned using Pittsburgh 30K (Pitts30K) training set [23]. Both the positive and negative samples are pre-computed before training to optimize the training duration. The training utilized the Adam optimizer, in conjunction with a cosine learning rate scheduler. The initial learning rate is set at $0.0005$. The training is continued until there is no further improvement in accuracy on the validation set.
@@
{tblr}
column2 = c, column3 = c, column4 = c, column6 = c, column7 = c, column8 = c, column14 = c, column15 = c, column16 = c, column18 = c, column19 = c, column20 = c, cell11 = r=2c, cell12 = c=3, cell15 = c, cell16 = c=3, cell110 = c=3c, cell114 = c=3, cell117 = c, cell118 = c=3, cell122 = c=3c,fg=black, cell25 = c, cell217 = c, cell222 = fg=black, cell223 = fg=black, cell224 = fg=black, cell35 = c, cell310 = c, cell311 = c, cell312 = c, cell317 = c, cell322 = c,fg=black, cell323 = c,fg=black, cell324 = c,fg=black, cell45 = c, cell410 = c, cell411 = c, cell412 = c, cell417 = c, cell422 = c,fg=black, cell423 = c,fg=black, cell424 = c,fg=black, cell51 = fg=black, cell52 = fg=black, cell53 = fg=black, cell54 = fg=black, cell56 = fg=black, cell57 = fg=black, cell58 = fg=black, cell510 = c,fg=black, cell511 = c,fg=black, cell512 = c,fg=black, cell514 = fg=black, cell515 = fg=black, cell516 = fg=black, cell518 = fg=black, cell519 = fg=black, cell520 = fg=black, cell522 = c,fg=black, cell523 = c,fg=black, cell524 = c,fg=black, cell61 = fg=black, cell62 = fg=black, cell63 = fg=black, cell64 = fg=black, cell65 = c, cell66 = fg=black, cell67 = fg=black, cell68 = fg=black, cell610 = c,fg=black, cell611 = c,fg=black, cell612 = c,fg=black, cell614 = fg=black, cell615 = fg=black, cell616 = fg=black, cell617 = c, cell618 = fg=black, cell619 = fg=black, cell620 = fg=black, cell622 = c,fg=black, cell623 = c,fg=black, cell624 = c,fg=black, cell75 = c, cell710 = c, cell711 = c, cell712 = c, cell717 = c, cell722 = c,fg=black, cell723 = c,fg=black, cell724 = c,fg=black, cell85 = c, cell810 = c, cell811 = c, cell812 = c, cell817 = c, cell822 = c,fg=black, cell823 = c,fg=black, cell824 = c,fg=black, cell910 = c, cell911 = c, cell912 = c, cell922 = c,fg=black, cell923 = c,fg=black, cell924 = c,fg=black, cell105 = c, cell1010 = c, cell1011 = c, cell1012 = c, cell1017 = c, cell1022 = c,fg=black, cell1023 = c,fg=black, cell1024 = c,fg=black, cell115 = c, cell1110 = c, cell1111 = c, cell1112 = c, cell1117 = c, cell1122 = c,fg=black, cell1123 = c,fg=black, cell1124 = c,fg=black, cell1222 = c,fg=black, cell1223 = c,fg=black, cell1224 = c,fg=black, cell1322 = c,fg=black, cell1323 = c,fg=black, cell1324 = c,fg=black, hline1,3,8,13-14 = -, hline2 = 2-16,18-24, Method & MSLS Validation [22] MSLS Challenge [22] Nordland [24] Pitts30K [23] Tokyo 24/7 [25] Robotcar-S2 [26]
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 .25m/2o .5m/5o 5.0m/10o
NetVLAD [2] 60.8 74.3 79.5 35.1 47.4 51.7 13.6 21.4 25.2 81.9 91.2 93.7 64.8 78.4 81.6 5.6 20.7 71.8
SFRS [27] 69.2 80.3 83.1 41.5 52.0 56.3 16.4 26.3 29.7 89.4 94.7 95.9 85.4 91.1 93.3 8.0 27.3 80.4
CosPlace [28] 85.2 92.3 93.2 60.9 71.7 76.7 54.7 70.9 77.9 89.0 94.7 96.1 81.0 90.8 93.7 8.2 29.9 83.7
MixVPR [29] 88.0 92.7 94.6 64.0 75.9 80.6 58.4 74.6 80.0 91.9 95.9 96.7 85.1 91.7 94.3 8.9 33.3 86.5
PlaceFormer (w/o re-ranking) 80.0 90.0 93.0 58.0 76.7 81.6 26.1 40.3 47.6 82.7 93.0 95.1 57.5 74.6 80.3 3.2 15.7 60.3
SP-SuperGlue [30] 78.1 81.9 84.3 50.6 56.9 58.3 29.1 33.5 34.3 87.2 94.8 96.4 88.2 90.2 90.2 9.5 35.4 85.4
DELG [12] 83.2 89.3 91.1 52.2 61.9 65.4 51.0 63.9 66.7 89.8 95.3 96.7 86.4 92.4 93.0 2.2 8.4 76.8
Patch-NetVLAD [5] 79.5 86.2 87.7 48.1 57.6 60.5 46.4 58.0 60.4 88.7 94.5 95.9 86.0 88.6 90.5 9.6 35.3 90.9
TransVPR [19] 86.8 91.2 92.4 63.9 74.0 77.5 58.8 75.0 78.7 89.0 94.9 96.2 79.0 82.2 85.1 9.8 34.7 80.0
$R^{2}$ormer [20] 89.7 95.0 96.2 73.0 85.9 88.8 60.6 66.8 68.7 91.1 95.2 96.3 88.6 91.4 91.7 10.5 35.2 85.2
PlaceFormer (Ours) 89.9 94.3 95.4 71.9 85.4 88.7 65.3 70.5 72.4 92.4 96.5 97.4 87.6 89.2 91.5 10.8 37.6 92.1
@@ IV-B Datasets
We evaluate PlaceFormer on multiple public benchmark datasets. These included the Nordland [24], Pittsburgh 30K (Pitts30K) [23], Tokyo24/7 [25], RobotCar Seasons v2 (RobotCar-S2) [26] and Mapillary Street-Level Sequences (MSLS) [22]. Each of these datasets offers a unique set of challenges, encompassing a wide array of environments and conditions that are crucial for thorough performance assessment. All images used in the evaluation are resized to a uniform resolution of $640\times 480$ to be consistent with other methods. The model trained on MSLS is used to evaluate MSLS and Nordland datasets. The model fine-tuned on Pitts30K is used to evaluate Pitts30K, RobotCar-S2, and Tokyo 24/7 (urban scenarios).
@@ IV-C Metrics
For MSLS, Nordland, and Pitta30K datasets, Recall@K is used as the primary metric for evaluation. This metric quantifies the percentage of query images that are correctly localized within a dataset. It does so by determining whether at least one of the top K-ranked reference images falls within a specified threshold distance from the query image. For our evaluation, we followed the precedent set in prior works [21, 19, 5], using a threshold distance of $25$ meters. The Recall@K is measured for K values of 1, 5, and 10. For the RobotCar-S2 dataset, we use the pose of the closest matching image as the estimated pose and compute recall under three default error thresholds.
@@ IV-D Comparison with State-of-the-arts
In our comparative analysis, PlaceFormer is benchmarked against a range of state-of-the-art methods to demonstrate its efficacy in visual place recognition. We evaluated it alongside methods such as NetVLAD [2], SFRS [27], CosPlace [28], and MixVPR [29], which primarily utilize global image representations. Alongside these methods, we also presented the results of our global retrieval to offer a comprehensive comparison. Furthermore, we compared PlaceFormer with techniques that employ both global and local features for retrieval and ranking. This included comparisons with Patch-NetVLAD [5] and DELG [12].
@@ V-A Quantitative Results
The quantitative performance of PlaceFormer, in comparison to other approaches, is detailed in Table I. PlaceFormer without local re-ranking outperforms traditional global retrieval methods like NetVLAD and SFRS in MSLS validation, MSLS challenge, and Nordland datasets, and yields comparable performance in Pitts30K and Tokyo 24/7 datasets. Furthermore, PlaceFormer without re-ranking outperforms multiple re-ranking methods as well, making it suitable to be used even without re-ranking based on specific requirements.
PlaceFormer with re-ranking achieves competitive results on all datasets. It outperforms all comparable methods on MSLS validation, Nordland, and Pitts30K in Recall@1 with absolute differences of $0.2\%$, $5.3\%$, and $1.3\%$ compared to the best-performing baseline of $R^{2}$ormer. When computing the average performance across all datasets, PlaceFormer demonstrates a substantial superiority over competing methods, surpassing NetVLAD, SFRS, CosPLace, MixVPR, SP-SuperGlue, DELG, PatchNetVLAD, TransVPR, and $R^{2}$ormer by margins of $30.8\%$, $21.04\%$, $11.6\%$, $3.94\%$, $14.78\%$, $8.9\%$, $11.68\%$, $5.92\%$, and $0.82\%$ in Recall@1 scores, respectively. While PlaceFormer generally surpasses other state-of-the-art methods, it falls short of outperforming $R^{2}$ormer on certain datasets such as MSLS Challenge and Tokyo 24/7. It is important to highlight that our method is specifically trained for global retrieval only, whereas $R^{2}$ormer undergoes training for re-ranking, contributing to its enhanced performance. PlaceFormer demonstrates superior performance compared to other methods across all three thresholds in the RobotCar S-2 dataset. Notably, it excels particularly under the $5.0m/10^{o}$ threshold, attributed to its adeptness in managing viewpoint discrepancies through multi-scale patch matching, distinguishing itself significantly from alternative approaches.